OpenStudy (anonymous):

3x^2-5x=1 <--- solve by the quadratic formula , set the equation to "0" first. must be in simplest radical form, show work please

OpenStudy (curry):

3x^2 - 5x -1 = 0

OpenStudy (anonymous):

\[x = \frac{5 \pm \sqrt{25 +12}}{6}\] \[x = \frac{5 \pm \sqrt{37}}{6} \]